OpenStudy (anonymous):

Ramona is playing a board game and rolls two number cubes. Let A = {the sum of the number cubes is odd} and let B = {the sum of the number cubes is divisible by 5}. List the outcomes in A ∩ B

OpenStudy (anonymous):

@jim_thompson5910 @jim_thompson5910 @beccaboo333 @Destinymasha @e.mccormick @NeedHelpQuick546 @ParthKohli @uri @whpalmer4 @zepdrix @zzr0ck3r @KingGeorge

OpenStudy (anonymous):

For A: The sum of two numbers is odd if and only if you're adding an even number to an odd one. For B: The sum is divisible by 5 if it ends in 5 or 0.

OpenStudy (anonymous):

@SithsAndGiggles So is it D?

OpenStudy (anonymous):

D? What is D?

OpenStudy (anonymous):

{1,3,5,7,9,11} {1,5} {3,5} {5}

OpenStudy (anonymous):

Strange... I'm not sure I understand the question. The way it's written: \[A=\left\{\text{sum of number cubes is odd}\right\}\] seems to suggest that the elements in \(A\) would be pairs of the form \((\star,\star)\) where each star is a component that is rolled (same for \(B\)). For example, \((5,2)\), whose sum is 7, would mean one cube rolled a 5 and the other rolled a 2. The means that if any element is shared among A and B, you'll have similar elements in the intersection, \((\star,\star)\). All of the answer choices are given as \((\star)\).

OpenStudy (anonymous):

So it is D

OpenStudy (anonymous):

It comes down to how you define A and B. My notion must be wrong. It could mean instead that A contains the possible sums that are odd, which would mean \[A=\{3,5,7,9,11\}\] and similarly, \[B=\{5,10\}\] The intersection would indeed be \(\{5\}\).

OpenStudy (anonymous):

It doesn't matter what I think, and it's not your fault, but I gotta say, this question is poorly written...

OpenStudy (anonymous):

@SithsAndGiggles Is this A?

OpenStudy (anonymous):

No - it's C

OpenStudy (anonymous):

What is this @SithsAndGiggles

OpenStudy (anonymous):

Can someone help with this? @aaronq @beccaboo333 @Destinymasha @e.mccormick @ranga @NeedHelpQuick546 @nelsonjedi

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Highest possible value of sin(x) is 1, and lowest is -1. so when you are asked to find the maximum and minimum, you just plug in 1 and -1 respectively.

OpenStudy (anonymous):

@geerky42 Still don't understand it

OpenStudy (anonymous):

Finding the min and max isn't necessary for this one - just the period. The period for a sine/cosine function of the form \(\sin bt\) is \(\dfrac{2\pi}{b}\). In this case, \(b=\dfrac{\pi}{5}\), so the period is \[\frac{2\pi}{\frac{\pi}{5}}=\frac{10\pi}{\pi}=10\]

OpenStudy (anonymous):

@SithsAndGiggles so the answer is which one

OpenStudy (anonymous):

The one that says the period is 10. There's only one.

OpenStudy (anonymous):

@SithsAndGiggles None of them have a period of 10

OpenStudy (anonymous):

@SithsAndGiggles The equation you gave me .4 so it has a period of 8 either A or D, which one

OpenStudy (anonymous):

You must be looking at a different question because I have no idea what you're talking about.

OpenStudy (anonymous):

Oh sorry yeah it is right

OpenStudy (pink.1d):

i hate